OpenStudy (anonymous):

One card is selected at random from a deck of cards. Determine the probability that the card is a 3 or 7.

OpenStudy (anonymous):

(1/52) + (1/52) = 2/52 = 1/26

OpenStudy (anonymous):

whoops...my mistake... there are 4 threes in a deck and 4 sevens 4/52 + 4/52 = 8/52 = 2/13